OpenStudy (anonymous):

Two identical jars are filled with equal numbers of marbles. The marbles are colored red or white. The ratio of red to white marbles is 7:1 in Jar I and 9:1 in Jar II. If there are 90 white marbles all together, how many red marbles are in Jar II?

OpenStudy (anonymous):

one way to say that the ratio of red to white is 7:1 is that \(\frac{7}{8}\) of the marbles in jar 1 are red, and similarly \(\frac{9}{10}\) of the marbles in jar 2 are red.

OpenStudy (anonymous):

\(\frac{1}{8}\) of the marbles in jar 1 are white and \(\frac{1}{10}\) of the marbles in jar two are white, and the total number of marbles in each jar are the same. we don't know how many are in each jar, so lets call in \(n\) is the number of marbles in each jar then you know \(\frac{1}{8}n+\frac{1}{10}n=90\)

OpenStudy (anonymous):

add and get \[\frac{9}{40}n=90\]so \[n=90\times \frac{40}{9}=400\]

OpenStudy (anonymous):

therefore each jar contains \(400\) marbles, and i bet you can finish it from there